OpenStudy (katey3324):

3: Suppose a line has slope-2/3 and passes through the point (1, 8). Which other point must also be on the graph? A) (7,4) B) (7,12) C) (-5,4) D) (-5, -12)

OpenStudy (katey3324):

b

sam (.sam.):

Nope

OpenStudy (katey3324):

c

sam (.sam.):

Wrong again

OpenStudy (katey3324):

d

sam (.sam.):

Have you did the equation first?

OpenStudy (katey3324):

i'm not good at math

OpenStudy (katey3324):

@logan13

sam (.sam.):

You should find the equation first then you could apply these coordinates A,B,C or D to find whether they fit in, you have a slope of, m=-2/3 and coordinate (1,8) to be \((x_1,y_1)\), using the standard \(y-y_1=m(x-x_1)\), your equation will be \(y=-\frac{2}{3}x+\frac{26}{3}\) Then, use the answers to see if they're correct by substituting in

OpenStudy (katey3324):

i'm not good at this

sam (.sam.):

Try using B (7,12), When x is 7, \[y=-\frac{2}{3}(7)+\frac{26}{3} \\ \\ y=-\frac{14}{3}+\frac{26}{3} \\ \\ y=12\] This means B, that point, is also on the equation of that graph

sam (.sam.):

oops its A, not B

sam (.sam.):

and y=4

sam (.sam.):

God my math is rusty